OpenStudy (anonymous):

Multiply. Write the result in the form a + bi 3i(2-3i)

OpenStudy (anonymous):

\(6i-9i^2\) is a start then since \(i^2=-1\) you get \(6i+9\) which in standard form is \(9+6i\)

OpenStudy (anonymous):

soo would (4 +5i) (2+i) be the same way.. I would multiply the two complex numbers together first right?

OpenStudy (anonymous):

no you can do this in one of two ways either multiply out using that loathsome "foil" i .e. do 4 multiplications, or just know that \[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\]

OpenStudy (anonymous):

so it would be 40i .. or am I just doing this whole thing wrong?

OpenStudy (anonymous):

i think probably you are

OpenStudy (anonymous):

\((8-5)+(4+10)i\) is a start

OpenStudy (anonymous):

where did you get those numbers?

OpenStudy (anonymous):

lets do it the slow way first, then i will show you where they came from

OpenStudy (anonymous):

oh okay. sorry about this.. i'm terrible at math..

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

suppose i had to multiply \[(4 +5x) (2+x) \] you would ge t \[8+4x+10x+5x^2=8+14x+5x^2\] clear or no?

OpenStudy (anonymous):

clear.

OpenStudy (anonymous):

ok lets repeat replacing \(x\) by \(i\) and go right to the answer \[8+14i+5i^2\] is that ok?

OpenStudy (anonymous):

let me know if it is not i replaced \(x\) by \(i\) and did the same thing as before, just didn't write down the steps this time, because they are identical

OpenStudy (anonymous):

yeah I got it

OpenStudy (anonymous):

ok one more step since \(i^2=-1\) you actually have \[8+14i-5\] ok with that?

OpenStudy (anonymous):

yes I am.

OpenStudy (anonymous):

and finally since \(8-5=3\) the "final answer" is \(3+14i\)

OpenStudy (anonymous):

now maybe it is clear where \[(8-5)+(4+10)i\] came from

OpenStudy (anonymous):

oops typo for \((a+bi)(c+di)\) when you multiply out you get \(bdi^2=-bd\) if this is confusing, you can always do it the way that we just did

OpenStudy (anonymous):

oh okaay thankyou!!!

OpenStudy (anonymous):

yw